A Solow Growth Model economy has a savings rate of 20 percent, a 5 percent rate of depreciation and a production function Y = K1/2L1/2. The economy currently has 100 workers and 400 units of capital. How much capital will the economy have one year from now?


1
Expert's answer
2021-01-30T06:31:28-0500
"Y=\sqrt K \sqrt L"

"L=100"

"Y=400"


"Y=200"

After 1 year


"400\times 1.05 \times 0.8=336"